OpenStudy (anonymous):

if s= 2x+t over r then x equals

OpenStudy (anonymous):

\[s=\frac{ 2x+t }{ r }\]is this what you mean:

OpenStudy (anonymous):

if so, it is very simple. multiply everything by r like so... \[r \left( s \right)=\left( \frac{ 2x+t }{ r } \right)r\] rs=2x+t, now subtract t from both sides rs-t= 2x, now divide by 2 on both sides \[x=\frac{ rs-t }{ 2 }\]